b. Which category has the fewest people? A. Category 4: People who are retired AND have grandchildren 0 B. Category 2: People who have grandchildren O C. Category 3: People who are retired OR have grandchildren 0 D. Category 1: People who are retired 0 E. There is not enough information to determine which category has the fewest people. H stand for heads and T stand for tails in an experiment where a fair coin is fiped twice. Assume outcomes listed are equally likely outcomes. ads and T stand for tails in an experiment where a fair coin is fipped twice. Assume that the four et HH, HT, TH, TT What are the probabilities of getting each of the following outcomes? a. Exactly 1 heacd b. At least 1 head c. No more than O heads Tos Total fips a. The probability of getting exactly 1 head is b. The probability of getting at least 1 head is2 c. The probability of getting no more than O heads is HH, HT, TH

I have the answers for number 9 but I need a better explanation and breaking it down.
Answer #1

a) P( exactly one head ) = No of Outcomes Favorable to Exactly one head/Total Outcomes

= 2/4

= 1/2

b) P( At least one head) = No of Outcomes Favorable to at least one head/Total Outcomes

= 3/4

c) P( No more than O head) = No of Outcomes Favorable to no more than 0 heads/Total Outcomes

= 1/4